OpenStudy (anonymous):

Find the slope. x/2+y/5=1

OpenStudy (precal):

multiply by 10 5x+2y=10 2y=10-5x y=5-(5/2)x

OpenStudy (anonymous):

thank you

OpenStudy (precal):

anytime :)